OpenStudy (kjfdkdklfd):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (kjfdkdklfd):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[\cos \theta = \frac{u\cdot v}{\|u\|\cdot\|v\|}\]

OpenStudy (unklerhaukus):

\[\|u\|=\sqrt{u_1^2+u_2^2}\]

OpenStudy (kjfdkdklfd):

Alright

OpenStudy (unklerhaukus):

so find \(\|u\|\), and \(\|v\|\) first

OpenStudy (kjfdkdklfd):

How I don't understand

OpenStudy (unklerhaukus):

The first vector is \[u = \langle2,-4\rangle\] so its length is \[\|u\|=\sqrt{(2)^2+(-4)^2}\]

OpenStudy (kjfdkdklfd):

So I have to solve that?

OpenStudy (unklerhaukus):

yeah simplify this into some irrational number

OpenStudy (kjfdkdklfd):

I just got square root of 2x^2 plus 16

OpenStudy (unklerhaukus):

\[\|u\|=\sqrt{(2)^2+(-4)^2}\\ \qquad=\sqrt{4+16}\\ \qquad=\]

OpenStudy (kjfdkdklfd):

2

OpenStudy (kjfdkdklfd):

Wait do you want the decimal or

OpenStudy (kjfdkdklfd):

2 times the square root of 5

OpenStudy (unklerhaukus):

keep it as an irrational number

OpenStudy (unklerhaukus):

yeah that's right ||u|| = 2√5

OpenStudy (unklerhaukus):

now find the length of the other vector ||v|| = √( 3^2 + (-8)^2 ) =

OpenStudy (kjfdkdklfd):

the square root of 73

OpenStudy (unklerhaukus):

right,

OpenStudy (unklerhaukus):

now what is the dot product of u and v?

OpenStudy (kjfdkdklfd):

Choices: 3.0° 6.0° -7.0° 16.0°

OpenStudy (unklerhaukus):

u * v = 2*3 + (-4)*(-8)

OpenStudy (kjfdkdklfd):

38

OpenStudy (unklerhaukus):

so we have \[\cos \theta = \frac{u\cdot v}{\|u\|\cdot\|v\|}\\ \qquad=\frac{38}{2\sqrt5\cdot\sqrt{73}}\]

OpenStudy (kjfdkdklfd):

19/square root of 365

OpenStudy (unklerhaukus):

the angle you are looking for is \[\theta = \arccos\left(\frac{19}{\sqrt{365}}\right)\] (make sure your calculator is set to degrees mode)

OpenStudy (kjfdkdklfd):

Is it 6.0 degrees

OpenStudy (kjfdkdklfd):

Can you just check my answer on this last question? I already did it I just want to make sure I'm right

OpenStudy (kjfdkdklfd):

Find a ⋅ b. a = <2, 4>, b = <2, 5> Choices: <4, 9> 24 <4, 20> 16

OpenStudy (unklerhaukus):

6 degrees is right,

OpenStudy (kjfdkdklfd):

I chose 24

OpenStudy (kjfdkdklfd):

But I also thought it might be <4,20>

OpenStudy (unklerhaukus):

for the dot product just remember to add the products of the corresponding components a ⋅ b = 2*2 + 4*5 = 4 + 20 = 24

OpenStudy (kjfdkdklfd):

Thanks!

OpenStudy (unklerhaukus):

bye!